OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 a. (–3, –1, –5) b. (–3, 5, –2) c. (2, –4, –3) d. (4, 4, –1)

OpenStudy (angela210793):

do u know how to use cramer?

OpenStudy (anonymous):

use substitution, or elimaination method and solve for 2 of the variables.

OpenStudy (campbell_st):

wasn't he that guy on seinfeld...?

OpenStudy (angela210793):

Me have no idea what seinfeld is O.o

OpenStudy (angela210793):

@bombshellbri do u know that or not? if u don't i can explain

OpenStudy (anonymous):

Or try adding the equations together in a way to make one variable disappear. Do this twice and then you have a 2x2 system.

OpenStudy (anonymous):

i dont not know

OpenStudy (anonymous):

x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 Add E2 + E3 to E1 5x + 3z = -30

OpenStudy (anonymous):

@angela210793 what is cramer?

OpenStudy (angela210793):

hmmm...well to use Cramer u must know matrices...do u know that? I'm insisting on Cramer cause it's very easy

OpenStudy (anonymous):

im sorry i dont know it either

OpenStudy (angela210793):

hmmm....well I guess @telliott99 can help u with his method...cause i make very masitakes in this method :/

OpenStudy (angela210793):

mistakes*

OpenStudy (anonymous):

alrite and thats okay

OpenStudy (anonymous):

Yes. The trouble with Cramer's rule for 3 x 3 is there is a lot of arithmetic.

OpenStudy (anonymous):

So we have one equation missing y. We need another one.

OpenStudy (anonymous):

x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 How about multiplying E2 by 3 and adding it to E1. I get 7x + 8z = -61 Before we had 5x + 3z = -30 Multiply the first one by -3 and the second one by 8 -21x -24z = 183 40x + 24z = -120 Add to get 19x = 63 That's pretty ugly. Maybe I made a mistake. Anyway, you see how to do that method. Maybe I'll try getting rid of z instead x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 Add twice E3 to both E1 and E2 5x - y = 14 6x -5y = -17 So y = 5x - 14 6x - 5(5x -14) = -17 -19x + 70 = -17 x = -87/-19 = 87/19 still ugly

OpenStudy (anonymous):

And one of those must have a mistake!

OpenStudy (anonymous):

Sorry I didn't solve your problem. 3 x 3 are hard enough, and harder with larger numbers and fractions, etc. But this and Cramer's rule are the two usual methods. http://en.wikipedia.org/wiki/Cramer's_rule

OpenStudy (anonymous):

Oh, I forgot.

OpenStudy (anonymous):

Since you have possible answers, you can always plug them into the equations.

OpenStudy (anonymous):

x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 a. (–3, –1, –5) b. (–3, 5, –2) c. (2, –4, –3) d. (4, 4, –1)

OpenStudy (anonymous):

Start with (d), they usually put the ones that work for most at the top.

OpenStudy (angela210793):

but he said he doesn't know matrices...that's why i quit and he can't do that Telliot....wht if he has 100 equations? is he really gonna try all the options O.o

OpenStudy (anonymous):

(d) does not work for E1

OpenStudy (anonymous):

im a girl* and does it matter which equation you use?

OpenStudy (anonymous):

No for 100 equations you get a computer (and the computer uses my method).

OpenStudy (anonymous):

We have eliminated (d) because it does not solve E1. Do you see that?

OpenStudy (angela210793):

sorry and that's not the right way to do it -_- she can't tell her teacher that she tried all the options to see which one satisifed all the equations :S

OpenStudy (anonymous):

sorry E3

OpenStudy (anonymous):

You wanna do Cramer's?

OpenStudy (anonymous):

If the method of solution is not specified, checking the answers seems perfectly valid to me.

OpenStudy (anonymous):

Particularly when they have not been taught methods of solution, apparently.

OpenStudy (anonymous):

i dont know what cramer's is. im only in algebra 2 ....

OpenStudy (anonymous):

Well, it's not too hard, if you want to try we can do it

OpenStudy (anonymous):

Actually, it will be a lot of work for a 3x3

OpenStudy (anonymous):

its an online class so we can try all the options to see which one satisifed all the equation

OpenStudy (anonymous):

Bingo

OpenStudy (angela210793):

I'm not telling to use cramer..there must be another way to solve this but whtever...do wht u want :P

OpenStudy (anonymous):

"must be"---there is not

OpenStudy (anonymous):

(b) does not solve E3

OpenStudy (anonymous):

Can you check the other two @bombshellbri

OpenStudy (anonymous):

hey telliot99 can you help me?

OpenStudy (anonymous):

using B?

OpenStudy (anonymous):

If I did it right, I found that (b) does not solve Eqn 3 and neither does (d) That leaves (a) and c as possible solutions.

OpenStudy (anonymous):

I think you can plug in the numbers to check?

OpenStudy (anonymous):

okay, ill try them

OpenStudy (anonymous):

Did you get an answer. I have one but I'm hoping you can tell me.

OpenStudy (anonymous):

(2x – y + 2z) - (x + 3y + 2z) = -15 + 16 x - 4y = 1 x = 1+4y (2x – 2y – z) - (2x – y + 2z) = 1 + 15 -y - 3z = 16 2(1+4y) - 2y - z = 2 + 6y - z = 1 6y - z = -1 18y - 3z = -3 -19y = 19, y = -1 x = -3 z = -5 answer: (A)

OpenStudy (anonymous):

No, sorry, we are no longer adding equations together x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 we are just trying this as a solution a. (–3, –1, –5) I get -3 + 3(-1) + 2(-5) = -16 check 2(-3) - (-1) + 2(-5) = -15 check 2(-3) -2(-1) -(-5) = 1 check

OpenStudy (anonymous):

Oh, I see you added them together correctly!

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

thank you very much :D

OpenStudy (anonymous):

yw